# Universal electrical transport of composite Fermi liquid to Metal transition in Moiré systems

### v1

@@ II.1 Critical Theory of QED-Chern-Simons coupled to spinon Fermi surface
For the CFL to FL transition at $\nu=\frac{1}{2}$, the charge-sector transition can be formulated as a bosonic Laughlin–superfluid transition at effective filling $\nu_{\Phi}=\frac{1}{2}$ (the electronic CFL corresponds to $\Phi$ in a Laughlin state, while the FL corresponds to $\Phi$ condensed). A convenient critical description of this bosonic transition uses a further fermionic partonization of the boson,
@@ III.1 Large-N expansion
$D_{a}(P)=\left(\frac{N}{2\pi}\epsilon_{\mu\nu\lambda}P_{\lambda}-\frac{1}{\epsilon}P_{\mu}P_{\nu}\right)^{-1}\equiv(N\boldsymbol{P}^{\prime})^{-1},$
@@
The polarization bubble $\Pi_{\psi}$ also yields the leading contribution to $\sigma_{\psi}$, which has been known in [26]. To zeroth order in $1/N$, one finds
@@ III.2 Low energy effective theory
$\left\{\begin{aligned} &\Delta_{a}^{-1}=-NP^{\prime}-N\Pi_{f}\\ &N\Pi_{b}=-\Delta_{b}^{-1}=N\Pi_{\psi}+N^{2}P\cdot\Delta_{a}\cdot P\end{aligned}\right$., (20)
@@
Solving the above equation set yields the tensor $\Pi_{b}$. To decouple the temporal component and spacial component of gauge field $b$, it’s much simpler to work under Coulumb gauge in which $\nabla\cdot\boldsymbol{b}=0$, and thus the spacial component of gauge field can be simplified to a scalar: $\boldsymbol{b}(\boldsymbol{q})=(\hat{\boldsymbol{z}}\times\hat{\boldsymbol{q}})b(\boldsymbol{q})$. The effective action turns into
@@ IV Critical Transport near the CFL to FL Transition
but the limiting regimes have distinct physical content: for $\hbar\omega\gg k_{B}T$ the response is dominated by phase-coherent motion of excitations created by the external field, while the DC limit $\hbar\omega\ll k_{B}T$ is governed by incoherent, inelastic scattering among thermally excited carriers and relaxation to local equilibrium. As shown explicitly for fractional quantum Hall critical points in [26], a calculation performed strictly at $T=0$ determines $\Sigma(\infty)$, whereas the DC conductivity is controlled by $\Sigma(0)$ and requires a finite-$T$ treatment of collisions. The appropriate framework in this hydrodynamic regime is a quantum kinetic equation: one first formulates the transport problem at $T>0$ and small $\omega$, and then takes the DC limit.
@@ IV.1 Quantum Boltzmann Equation
$\psi_{1}(x,t)=\int_{\boldsymbol{k}}\left[\gamma_{+}(\boldsymbol{k},t)\begin{pmatrix}\frac{1}{\sqrt{2}}\\ \frac{\tilde{k}}{\sqrt{2}}\end{pmatrix}+\gamma_{-}(\boldsymbol{k},t)\begin{pmatrix}\frac{1}{\sqrt{2}}\\ -\frac{\tilde{k}}{\sqrt{2}}\end{pmatrix}\right]e^{i\boldsymbol{k}\cdot\boldsymbol{x}}$ (27)
@@ V.1 Effective propagator
Utilizing the same method as which in Sec.III.2, we can write down the propagator equations by order:
$\left\{\begin{aligned} &\Delta_{b_{2}}^{-1}=-NP^{\prime}-N\Pi_{\psi},\quad\Delta_{a}^{-1}=-N\Pi_{f}\\ &\Delta_{\alpha}^{-1}=-2NP^{\prime}-N^{2}P\cdot\Delta_{b_{2}}\cdot P-N^{2}P\cdot\Delta_{a}\cdot P\\ &N\Pi_{b}=-\Delta_{b_{1}}^{-1}=NP^{\prime}+N\Pi_{\psi}+N^{2}P\cdot\Delta_{\alpha}\cdot P\end{aligned}\right$. (47)
@@ VI.1 Conductivity at large frequencies
$\sigma_{\psi}(\omega\gg T)\approx\frac{N}{16}\frac{e^{2}}{h}.$ (49)
The full response function $\sigma(\omega/T)$ thus encodes a universal crossover between these two limits. Given that the physical resistivity is the sum of the critical sector $\rho_{\psi}$ and the spectator Fermi surface $\rho_{f}$, the critical sector acts as a robust series resistor. This ensures that the crossover in $\sigma_{\psi}$ should be observable as a frequency-dependent modulation of the total resistivity $\rho_{xx}$.
@@ Appendix A Analytical treatments of gauge polarization function
And we have $\boldsymbol{P}=\frac{P}{2\pi}P_{O}$, $\boldsymbol{P}^{\prime}=\frac{P}{2\pi}P_{O}-\frac{1}{\epsilon}E$. We can extrat these components via
@@ A.2 $\nu=-3/4$ case
$\Delta_{b}^{-1}=-\boldsymbol{P}^{\prime}-\Pi_{\psi}-\boldsymbol{P}\cdot\Delta_{\alpha}\cdot\boldsymbol{P},$ (65)
$\Delta_{a}^{-1}=-\Pi_{f}-\boldsymbol{P}\cdot\Delta_{\alpha}\cdot\boldsymbol{P},$ (66)
$\Delta_{a}^{-1}=-2\boldsymbol{P}^{\prime}-\boldsymbol{P}\cdot\Delta_{b_{1}}\cdot\boldsymbol{P}-\boldsymbol{P}\cdot\Delta_{b_{2}}\cdot\boldsymbol{P}-\boldsymbol{P}\cdot\Delta_{a}\cdot\boldsymbol{P},$ (67)
$\Delta_{A}^{-1}=-\boldsymbol{P}\cdot\Delta_{\alpha}\cdot\boldsymbol{P},$ (68)

### v2

@@ II.1 Critical Theory of QED-Chern-Simons coupled to spinon Fermi surface
For the CFL to FL transition at $\nu=-\frac{1}{2}$, the charge-sector transition can be formulated as a bosonic Laughlin–superfluid transition at effective filling $\nu_{\Phi}=\frac{1}{2}$ (the electronic CFL corresponds to $\Phi$ in a Laughlin state, while the FL corresponds to $\Phi$ condensed). A convenient critical description of this bosonic transition uses a further fermionic partonization of the boson,
@@ III.1 Large-N expansion
$D_{a}(P)=\left(\frac{N}{2\pi}\epsilon_{\mu\nu\lambda}P_{\lambda}-\frac{1}{\epsilon}P_{\mu}P_{\nu}\right)^{-1}\equiv(N\boldsymbol{\Theta}^{\prime})^{-1},$
@@
The polarization bubble $\Pi_{\psi}$ also yields the leading contribution to $\sigma_{\psi}$, which has been known in [26]. To zeroth order in $1/N$, one finds $\sigma_{\psi}$ can be divided into the thermally excited quasi-particle part($\sigma_{\mathrm{qp}}$), and the contribution from the creation of quasiparticle-quasihole pairs by the external sources($\sigma_{\mathrm{coh}}$):
@@ III.2 Low energy effective theory
$\left\{\begin{aligned} &\Delta_{a}^{-1}=-N\boldsymbol{\Theta}^{\prime}-N\Pi_{f}\\ &N\Pi_{b}=-\Delta_{b}^{-1}=N\Pi_{\psi}+N^{2}\boldsymbol{\Theta}\cdot\Delta_{a}\cdot\boldsymbol{\Theta}\end{aligned}\right$. (20)
@@
Solving the above equation set yields the tensor $\Pi_{b}$. To decouple the temporal component and spacial component of gauge field $b$, it is much simpler to work under Coulumb gauge in which $\nabla\cdot\boldsymbol{b}=0$, and thus the spacial component of gauge field can be simplified to a scalar: $\boldsymbol{b}(\boldsymbol{q})=(\hat{\boldsymbol{z}}\times\hat{\boldsymbol{q}})b(\boldsymbol{q})$. The effective action turns into
@@ IV Critical Transport near the CFL to FL Transition
where the limiting regimes have distinct physical content: for $\hbar\omega\gg k_{B}T$ the response is dominated by phase-coherent motion of excitations created by the external field, while the DC limit $\hbar\omega\ll k_{B}T$ is governed by incoherent, inelastic scattering among thermally excited carriers and relaxation to local equilibrium. As shown explicitly for fractional quantum Hall critical points in [26], a calculation performed strictly at $T=0$ determines $\Sigma(\infty)$, whereas the DC conductivity is controlled by $\Sigma(0)$ and requires a finite-$T$ treatment of collisions. The appropriate framework in this hydrodynamic regime is a quantum kinetic equation: one first formulates the transport problem at $T>0$ and small $\omega$, and then takes the DC limit.
@@ IV.1 Quantum Boltzmann Equation
$\psi_{1}(x,t)=\int_{\boldsymbol{k}}\left[\gamma_{+}(\boldsymbol{k},t)\begin{pmatrix}\frac{1}{\sqrt{2}}\\ \frac{\tilde{k}}{\sqrt{2}}\end{pmatrix}+\gamma_{-}(\boldsymbol{k},t)\begin{pmatrix}\frac{1}{\sqrt{2}}\\ -\frac{\tilde{k}}{\sqrt{2}}\end{pmatrix}\right]e^{i\boldsymbol{k}\cdot\boldsymbol{x}},$ (27)
@@ V.1 Effective propagator
Utilizing the same method as in Sec.III.2, we can write down the propagator equations by order:
$\left\{\begin{aligned} &\Delta_{b_{2}}^{-1}=-N\boldsymbol{\Theta}^{\prime}-N\Pi_{\psi},\quad\Delta_{a}^{-1}=-N\Pi_{f}\\ &\Delta_{\alpha}^{-1}=-2N\boldsymbol{\Theta}^{\prime}-N^{2}\boldsymbol{\Theta}\cdot\Delta_{b_{2}}\cdot\boldsymbol{\Theta}-N^{2}\boldsymbol{\Theta}\cdot\Delta_{a}\cdot\boldsymbol{\Theta}\\ &N\Pi_{b}=-\Delta_{b_{1}}^{-1}=N\boldsymbol{\Theta}^{\prime}+N\Pi_{\psi}+N^{2}\boldsymbol{\Theta}\cdot\Delta_{\alpha}\cdot\boldsymbol{\Theta}\end{aligned}\right$. (47)
@@ VI.1 Conductivity at large frequencies
$\sigma_{\psi}(\omega\gg T)\approx\frac{N}{16}\frac{e^{2}}{h},$ (49)
which can also be read from Eq.18. The full response function $\sigma(\omega/T)$ thus encodes a universal crossover between these two limits. Given that the physical resistivity is the sum of the critical sector $\rho_{\psi}$ and the spectator Fermi surface $\rho_{f}$, the critical sector acts as a robust series resistor. This ensures that the crossover in $\sigma_{\psi}$ should be observable as a frequency-dependent modulation of the total resistivity $\rho_{xx}$.
@@ Appendix A Analytical treatments of gauge polarization function
And we have $\boldsymbol{\Theta}=\frac{P}{2\pi}P_{O}$, $\boldsymbol{\Theta}^{\prime}=\frac{P}{2\pi}P_{O}-\frac{1}{0^{+}}E$. We can extract these components via
@@ A.2 $\nu=-3/4$ case
$\Delta_{b}^{-1}=-\boldsymbol{\Theta}^{\prime}-\Pi_{\psi}-\boldsymbol{\Theta}\cdot\Delta_{\alpha}\cdot\boldsymbol{\Theta},$ (65)
$\Delta_{a}^{-1}=-\Pi_{f}-\boldsymbol{\Theta}\cdot\Delta_{\alpha}\cdot\boldsymbol{\Theta},$ (66)
$\Delta_{a}^{-1}=-2\boldsymbol{\Theta}^{\prime}-\boldsymbol{\Theta}\cdot\Delta_{b_{1}}\cdot\boldsymbol{\Theta}-\boldsymbol{\Theta}\cdot\Delta_{b_{2}}\cdot\boldsymbol{\Theta}-\boldsymbol{\Theta}\cdot\Delta_{a}\cdot\boldsymbol{\Theta},$ (67)
$\Delta_{A}^{-1}=-\boldsymbol{\Theta}\cdot\Delta_{\alpha}\cdot\boldsymbol{\Theta},$ (68)